# Tunneling dynamics of ${}^{164}$y supersolids and droplets

### v1

@@ Abstract
The tunneling dynamics of a magnetic ${}^{164}$y quantum gas in an elongated or circular skewed double-well trap is investigated with a time-dependent extended Gross-Pitaevskii approach. Upon lifting the energy offset, different tunneling regimes can be identified. In the elongated trap and for sufficiently large offset, the different configurations exhibit collective macroscopic tunneling. For smaller offset, partial reflection from and transmission through the barrier lead to density accumulation in both wells, and eventually to tunneling-locking. One can also reach the macroscopic self-trapping regime for increasing relative dipolar interaction strength, while tunneling vanishes for large barrier heights. A richer dynamical behavior is observed for the circular trap. For instance, the supersolid maintains its shape, while the superfluid density gets distorted signifying the emergence of peculiar excitation patterns in the macroscopic tunneling regime. The findings reported here may offer new ways to probe distinctive dynamical features in the supersolid and droplet regimes.
@@ I Introduction
Advances in ultra-cold atoms have ushered new possibilities to observe this intricate SS state of matter. Predictions were made for Rydberg-excited BECs Henkel et al. (2010) and in the dipole-blockade regime Cinti et al. (2010). Early experimental realizations of supersolidity were reported for example in spin-orbit Li et al. (2017) and BECs coupled to optical cavities Léonard et al. (2017a, b). The above mentioned dBECs, however, provide a particularly favorable setting since the localization is not externally initiated, but rather is an intrinsic property resulting from the spontaneous symmetry breaking due to the interactions. Only shortly after the discovery of dipolar droplets, the SS state was uncovered in three milestone experiments with highly magnetic dysprosium and erbium atoms Böttcher et al. (2019); Tanzi et al. (2019a); Chomaz et al. (2019).
In dBECs at the verge to SS and self-bound states, quantum fluctuation contributions play a major role. Similar to the description of stable self-bound droplets in binary BECs Petrov (2015), they are commonly modelled in the extended Gross-Pitaevskii equation (eGPE) Wächtler and Santos (2016a, b); Bisset et al. (2016); Chomaz et al. (2016); Baillie and Blakie (2018) including the Lee-Huang-Yang (LHY) correction term Lee et al. (1957); Lima and Pelster (2011) to approximate quantum fluctuations to first order. In a dipolar SS, the spontaneous breaking of translational symmetry Böttcher et al. (2019); Tanzi et al. (2019a); Chomaz et al. (2019); Norcia et al. (2021) gives rise to density modulations while partially retaining superfluid (SF) properties and exhibits both diagonal and off-diagonal long-range order Gross (1957, 1958); Yang (1962); Andreev and Lifshits (1969); Chester (1970); Leggett (1970). This phenomenon is underpinned by the presence of a roton minimum in the energy-momentum dispersion relation which has been experimentally probed Chomaz et al. (2018); Guo et al. (2019); Natale et al. (2019). Notably, in cases where long-range interactions are dominant, the dilute SF background density, characteristic of the SS phase, may vanish leading to the formation of isolated droplet lattice (DL) configurations Ferrier-Barbut et al. (2016); Böttcher et al. (2019); Chomaz et al. (2016).
@@
For BECs with short-range interactions, the tunneling dynamics in double-well potentials has been widely investigated, see for instance the review Gati and Oberthaler (2007) or the collected articles in Keshavamurthy and Schlagheck (2011). Such systems constitute atomic analogues of the well-known superconducting Josephson junction Josephson (1962) and exhibit tunneling phenomena Albiez et al. (2005); Gati and Oberthaler (2007) including, for instance, Josephson oscillations Smerzi et al. (1997) and macroscopic quantum self-trapping Raghavan et al. (1999). However, the tunneling dynamics of dBECs in the SS and DL phases and in particular, the interplay of short- and long-range interactions on the emergent tunneling regimes still remain largely unexplored.
@@
Here, we report an attempt to theoretically investigate the tunneling dynamics of a dipolar SS or DL in a double-well potential in the eGPE framework. The approach is anticipated to capture the basic tunneling regimes of the dBEC, whilst higher-order effects, such as additional tunneling channels or inter-band contributions, require more sophisticated approaches Mistakidis et al. (2023) and are left for future work.
@@
The dBEC is initially confined in a tilted double-well potential, with the individual wells chosen to be of elongated or circular shape. Then system is suddenly quenched from an initial energy offset between the wells to a symmetric double-well potential. The initial energy offset, the relative dipolar interaction strength (i.e., contact versus dipolar coupling strength) and the barrier height determine the tunneling characteristics of the system. Prior to the quench, the single-well configurations exhibit the well-known formation of SF, SS, and DL states (see e.g. the review of Ref. Chomaz et al. (2022)). For an elongated double-well with a sufficiently large initial energy offset, the different dBEC configurations exhibit periodic oscillations between the wells with constant amplitude and frequency, thus featuring collective tunneling. When the offset is reduced, partial reflection and transmission occur through the barrier, resulting in the accumulation of density in both wells and eventually leading to tunneling-locking. At long evolution times, in the self-bound DL regime, a significant population difference is established, distinguishing it from the SF phase where the imbalance vanishes. Additionally, the SS and DL states exhibit a reduced center-of-mass velocity compared to the SF phase, indicating their rigidity. Increasing the barrier height while keeping all other parameters fixed, it is possible to enter the macroscopic self-trapping regime. The different patterns also manifest in systems with circular single-well traps, however with a richer phenomenology.
@@
This work proceeds as follows: Section II introduces the dBEC setting and the relevant eGPE framework. The ground-state phases of the system confined in a double-well with an energy offset for varying interactions are discussed in Sec. III. Consecutively, the resultant tunneling dynamics of the different phases after suddenly switching off the underlying energy offset is analyzed for a double-well with elongated single traps in Sec. IV and with circular ones in Sec. V. We summarize and comment on future perspectives in Sec. VI. In Appendix A, we discuss the effect of three-body recombination on the tunneling behavior, while Appendix B explicates the persistence of tunneling after a linear ramp of the energy offset. Appendix C further elaborates on some technical details.
@@ II Extended Gross-Pitaevskii for the dipolar gas
We consider a dBEC of ${}^{164}$y atoms having a magnetic dipole moment $\mu_{m}=9.93\mu_{B}$ (where $\mu_{B}$ is the Bohr magneton) being polarized along the $z$-direction. At zero temperature the dynamics of the dBEC is modeled by the eGPE Wächtler and Santos (2016a, b); Bisset et al. (2016); Chomaz et al. (2016); Baillie and Blakie (2018) containing the first order LHY beyond mean-field correction,
@@
The three-dimensional (3D) wave function is represented by $\psi(\textbf{r},t)$ and $m$ is the atom mass. The long-range anisotropic DDI is described by the dipolar interaction $U_{dd}(\textbf{r},t)=\frac{\mu_{0}\mu^{2}_{m}}{4\pi}\left[\frac{1-3\cos^{2}\theta}{\mathbf{r}^{3}}\right]$. The angle between the relative distance $\mathbf{r}$ of two dipoles and the $z$ axis of the magnetic field is denoted by $\theta$, and $\mu_{0}$ represents the permeability of the vacuum. Moreover, the dipolar atoms experience short-range contact interactions whose strength is quantified via the 3D $s$-wave scattering length, $a_{s}$, that can be experimentally tuned through Fano-Feshbach resonances Koch et al. (2008); Ilzhöfer et al. (2021). The LHY term Lee et al. (1957) approximates the lowest-order quantum fluctuation contributions to the energy functional, which in 3D are repulsive and scale with the gas density as $\sim n^{3/2}$. It takes the form $\gamma(\epsilon_{{\rm dd}})=\frac{32}{3}g\sqrt{\frac{a_{s}^{3}}{\pi}}\left(1+\frac{3}{2}\epsilon_{{\rm dd}}^{2}\right)$ Lima and Pelster (2011). Its contribution is crucial for the system to sustain many-body self-bound states such as droplets and SS and it has been shown to adequately describe experimental observations (as reviewed in Chomaz et al. (2022)).
In a 3D harmonic trap the relative strength $\epsilon_{{\rm dd}}=a_{{\rm dd}}/a_{s}$ between the DDI and the short-range interactions determines the many-body phase of the system Ferrier-Barbut et al. (2016). In particular, the dipolar length of ${}^{164}$y atoms is $a_{{\rm dd}}=\mu_{0}\mu^{2}_{m}m/12\pi\hbar^{2}=131a_{B}$, with $a_{B}$ being the Bohr radius. For sufficiently small values of $\epsilon_{{\rm dd}}$, the system exhibits a SF phase. As the scattering length decreases and $\epsilon_{{\rm dd}}$ increases, the attractive component of the dipolar force and the LHY term begin to play a more significant role in defining the equilibrium solution. Within a small range of $\epsilon_{\rm dd}$, the system favors the SS phase, where a periodic structure of localized densities coherently connected by a dilute superfluid background emerges. However, a further increase in $\epsilon_{{\rm dd}}$ leads to a DL phase.
The initial state is prepared in an external 3D tilted double-well potential that reads
@@
where $V_{D}$ models the height of the central double-well barrier and $w$ is its width. For our purposes, we use $V_{D}=10\hbar\omega_{x}$ and $w=0.5l_{{\rm osc}}$, with $l_{{\rm osc}}=\sqrt{\hbar/(m\omega_{x})}$ being the harmonic oscillator length and $\omega_{x}$ denoting the frequency of the external confinement. The last term is the linear external field gradient of strength $D$ that energetically favors the population of the left well. The tilt strength ranges from $D=1\hbar\omega_{x}/l_{{\rm osc}}$ to $D=10\hbar\omega_{x}/l_{{\rm osc}}$.
@@
Our analysis extends from (i) an elongated double-well characterized by $(\omega_{x},\omega_{y},\omega_{z})=2\pi\times(19,53,81)\rm Hz$ (with the single wells being similar to those in Tanzi et al. (2019b); Böttcher et al. (2019)) to (ii) a double-well with circular single wells, $(\omega_{x},\omega_{y},\omega_{z})=2\pi\times(43,43,133)\rm Hz$ (as in Klaus et al. (2022)). The lowest panels of Fig. 1 show the contours of the potentials cut through $z=0$ for $D=10\hbar\omega_{x}/l_{\rm osc}$. Also, the upper panels (a)-(c) and (d)-(f) represent the dBEC configurations in the left well before the quench for the elongated and planar single wells respectively. The characteristic timescales determined by the trap frequency accordingly correspond to $\omega_{x}^{-1}=8.4$ ms and $\omega_{x}^{-1}=3.7$ ms in the elongated and planar cases, with harmonic oscillator lengths $l_{{\rm osc}}=\sqrt{\hbar/(m\omega_{x})}=1.8{\rm\mu m}$ and $1.2{\rm\mu m}$, respectively. Another important timescale is set by the energy difference among the pre- and post-quench states, $\tau_{D_{{\rm in}}}=1/\Delta\mathcal{E}_{D_{\rm in}}$, where $\Delta\mathcal{E}_{D_{\rm in}}\equiv\mathcal{E}_{D_{{\rm in}}}-\mathcal{E}_{{\rm fi}}$. Here, $\mathcal{E}_{D_{{\rm in}}}$ and $\mathcal{E}_{D_{{\rm fi}}}$ refer to the energy of the dBEC at the initial ($D_{{\rm in}}=D$) and final ($D_{{\mathrm{f}}}=0$) tilt strength. The interaction dependence of $\Delta\mathcal{E}_{D_{\rm in}}$ is weak for the considered values of $\epsilon_{{\rm dd}}$. We thus provide an average value for each $D_{{\rm in}}$. In this sense, relevant timescales obtained from this energy difference correspond to $\tau_{D_{{\rm in}}=10}\approx 40$ms and $\tau_{D_{{\rm in}}=7}\approx 20$ms in both settings.
@@
In what follows, we explore a ${}^{164}$y dBEC under the influence of the energy offset for specific values of $\epsilon_{{\rm dd}}$ corresponding to a SF, a SS, and a DL in the initial state. After preparing the initial configuration, we perform a quench of the tilt strength from a finite value to zero, and monitor the system dynamics as the energy offset vanishes for up to one second.
@@
Figure 1: Ground state density iso-surfaces of a dBEC initially localized on the left side of a tilted double-well trap with (a)-(c) elongated and (d)-(f) circular single-well shapes, as depicted by the contour lines in the lowest panels. The density isosurfaces represent 20% and 2.5% of maximum density. Depending on the relative interaction strength $\epsilon_{{\rm dd}}$ the dBEC in the left well features (a), (d) a regular SF for $\epsilon_{dd}=1.31$, (b) an elongated ($\epsilon_{{\rm dd}}=1.4$) or (e) planar SS ($\epsilon_{{\rm dd}}=1.49$), (c) an elongated DL pattern for $\epsilon_{dd}=1.49$ as well as (f) a single droplet surrounded by a homogeneous density distribution for $\epsilon_{{\rm dd}}=1.559$. Notice that the individual transitions are shifted to larger $\epsilon_{{\rm dd}}$ from the quasi-1D to the quasi-2D setting. The tilt strength is $D=7\hbar\omega_{x}/l_{{\rm osc}}$ enforcing all $N=40000$ magnetic atoms to reside on the left part of the double-well. The latter is characterized by frequencies (a)-(c) $(\omega_{x},\omega_{y},\omega_{z})=2\pi\times(19,53,81)\rm Hz$ and (d)-(f) $(\omega_{x},\omega_{y},\omega_{z})=2\pi\times(43,43,131)\rm Hz$, while having a width $w=0.5l_{{\rm osc}}$ and height $V_{D}=10\hbar\omega_{x}$. The potential contours at $z=0$ plane for $D=7\hbar\omega_{x}/l_{{\rm osc}}$, along with the variation of the potentials $V(x,0,0)$ along $x$ for $D=7\hbar\omega_{x}/l_{{\rm osc}}$ (solid line) and $D=0$ (dashed line), are shown in the left and right panels for elongated and circular single-wells, respectively.

### v2

# Tunneling dynamics of 164Dy supersolids and droplets

@@ Abstract
The tunneling dynamics of a magnetic 164Dy quantum gas in an elongated or pancake skewed double-well trap is investigated with a time-dependent extended Gross-Pitaevskii approach. Upon lifting the energy offset, different tunneling regimes can be identified. In the elongated trap and for sufficiently large offset, the different configurations exhibit collective macroscopic tunneling. For smaller offset, partial reflection from and transmission through the barrier lead to density accumulation in both wells, and eventually to tunneling-locking. One can also reach the macroscopic self-trapping regime for increasing relative dipolar interaction strength, while tunneling vanishes for large barrier heights. A richer dynamical behavior is observed for the pancake-like trap. For instance, the supersolid maintains its shape, while the superfluid density gets distorted signifying the emergence of peculiar excitation patterns in the macroscopic tunneling regime. The findings reported here may offer new ways to probe distinctive dynamical features in the supersolid and droplet regimes.
@@ I Introduction
Advances in ultra-cold atoms have opened up new possibilities to observe this intricate SS state of matter. Predictions were made for Rydberg-excited BECs Henkel et al. (2010) and in the dipole-blockade regime Cinti et al. (2010). Early experimental realizations of supersolidity were reported for example in spin-orbit Li et al. (2017) and BECs coupled to optical cavities Léonard et al. (2017a, b). The above mentioned dBECs, however, provide a particularly favorable setting since the localization is not externally initiated, but rather is an intrinsic property resulting from the spontaneous symmetry breaking due to the interactions. Only shortly after the discovery of dipolar droplets, the SS state was uncovered in three milestone experiments with highly magnetic dysprosium and erbium atoms Böttcher et al. (2019); Tanzi et al. (2019a); Chomaz et al. (2019).
In dBECs, for the SS and self-bound states, quantum fluctuation plays a major role. Similar to the description of stable self-bound droplets in binary BECs Petrov (2015), they are commonly modelled with the extended Gross-Pitaevskii method (eGPE) Wächtler and Santos (2016a, b); Bisset et al. (2016); Chomaz et al. (2016); Baillie and Blakie (2018), including the Lee-Huang-Yang (LHY) correction Lee et al. (1957); Lima and Pelster (2011) to approximate quantum fluctuations to first order. In a dipolar SS, the spontaneous breaking of translational symmetry Böttcher et al. (2019); Tanzi et al. (2019a); Chomaz et al. (2019); Norcia et al. (2021) gives rise to density modulations while partially retaining superfluid (SF) properties and exhibits both diagonal and off-diagonal long-range order Gross (1957, 1958); Yang (1962); Andreev and Lifshits (1969); Chester (1970); Leggett (1970). This phenomenon is underpinned by the presence of a roton minimum in the energy-momentum dispersion relation which has been experimentally probed Chomaz et al. (2018); Guo et al. (2019); Natale et al. (2019). Notably, in cases where long-range interactions are dominant, the dilute SF background density, characteristic of the SS phase, may vanish leading to the formation of isolated droplet lattice (DL) configurations Ferrier-Barbut et al. (2016); Böttcher et al. (2019); Chomaz et al. (2016).
@@
For BECs with short-range interactions, the tunneling dynamics in double-well potentials has been widely investigated, see for instance the review Gati and Oberthaler (2007) or the collected articles in Keshavamurthy and Schlagheck (2011). Such systems constitute atomic analogues of the well-known superconducting Josephson junction Josephson (1962) and exhibit tunneling phenomena Albiez et al. (2005); Gati and Oberthaler (2007) including, for instance, Josephson oscillations Smerzi et al. (1997) and macroscopic quantum self-trapping Raghavan et al. (1999). However, the tunneling dynamics of dBECs in the SS and DL phases and in particular, the interplay of short- and long-range interactions on the emergent tunneling regimes remain largely unexplored. Additionally, since the droplet crystal configuration depends on the spatial dimension of the system, it would be intriguing to explore the impact of dimensionality on the emergent tunneling behavior. Relevant open questions, for instance, whether the crystal arrangement retains its symmetry in the course of tunneling and how suppression of the latter occurs due to interactions in various dimensions, are addressed.
@@
Here, we report an attempt to theoretically investigate the tunneling dynamics of a dipolar SS or DL in a double-well potential in the eGPE framework. The approach is anticipated to capture the basic dBEC tunneling dynamics, whilst higher-order effects, such as additional tunneling channels or inter-band contributions, require more sophisticated approaches Mistakidis et al. (2023); Wilsmann et al. (2018) and are left for future work.
@@
The dBEC is initially confined in a tilted double-well potential, with the individual wells chosen to be elongated or pancake-like. Then the system is suddenly quenched from an initial energy offset between the wells to a symmetric double-well potential. The initial energy offset, the relative dipolar strength (i.e., contact versus dipolar coupling) and the barrier height determine the tunneling characteristics of the system. Prior to the quench, the single-well configurations exhibit the well-known formation of SF, SS, and DL states (see e.g. the review of Ref. Chomaz et al. (2022)). For an elongated double-well with a sufficiently large initial energy offset, the different dBEC configurations exhibit periodic oscillations between the wells with constant amplitude and frequency, thus featuring collective tunneling. When the offset is reduced, partial reflection and transmission occur through the barrier, resulting in the accumulation of density in both wells and eventually leading to tunneling-locking. At long evolution times, in the self-bound DL regime, a significant population difference is established, distinguishing it from the SF phase where the imbalance vanishes. Additionally, the SS and DL states exhibit a reduced center-of-mass velocity compared to the SF phase, indicating their rigidity. Increasing the barrier height while keeping all other parameters fixed, it is possible to enter the macroscopic self-trapping regime. The different patterns also manifest in pancake-like double-well systems, however with a richer phenomenology.
@@
This paper proceeds as follows: Section II introduces the dBEC setting and the relevant eGPE framework. The ground-state phases of the system confined in a double-well with an energy offset for varying interactions are discussed in Sec. III. Consecutively, the resultant tunneling dynamics of the different phases after suddenly switching off the underlying energy offset is analyzed for an elongated double-well in Sec. IV and pancake-like one in Sec. V. We summarize and comment on future perspectives in Sec. VI. In Appendix A, we discuss the effect of three-body recombination on the tunneling behavior, while Appendix B explicates the persistence of tunneling after a linear ramp of the energy offset. Appendix C further elaborates on some technical details.
@@ II Extended Gross-Pitaevskii for the dipolar gas
We consider a dBEC of 164Dy atoms having a magnetic dipole moment $\mu_{m}=9.93\mu_{B}$ (where $\mu_{B}$ is the Bohr magneton) being polarized along the $z$-direction. At zero temperature, the dynamics of the dBEC is modeled by the eGPE Wächtler and Santos (2016a, b); Bisset et al. (2016); Chomaz et al. (2016); Baillie and Blakie (2018) containing the first order LHY beyond mean-field correction,
@@
The three-dimensional (3D) wave function is represented by $\psi(\textbf{r},t)$ and $m$ is the mass of Dy atom. The long-range anisotropic DDI is $U_{dd}(\textbf{r},t)=\frac{\mu_{0}\mu^{2}_{m}}{4\pi}\left[\frac{1-3\cos^{2}\theta}{\mathbf{r}^{3}}\right]$. The angle between the relative distance $\mathbf{r}$ of two dipoles and the $z$ axis of the quantization axis (defined by the magnetic field) is denoted by $\theta$, and $\mu_{0}$ represents the permeability of the vacuum. Moreover, the dipolar atoms experience short-range contact interactions whose strength is quantified via the 3D $s$-wave scattering length, $a_{s}$, that can be experimentally tuned through Fano-Feshbach resonances Koch et al. (2008); Ilzhöfer et al. (2021). The LHY term Lee et al. (1957) approximates the lowest-order quantum fluctuation contributions to the energy functional. These fluctuations in 3D appear to be repulsive and scale with the gas density as $\sim n^{3/2}$, taking the form $\gamma(\epsilon_{{\rm dd}})=\frac{32}{3}g\sqrt{\frac{a_{s}^{3}}{\pi}}\left(1+\frac{3}{2}\epsilon_{{\rm dd}}^{2}\right)$ Lima and Pelster (2011). Its contribution is crucial for the system to sustain many-body self-bound states such as droplets and SS and it has been shown to adequately describe experimental observations (as reviewed in Chomaz et al. (2022)).
In a 3D harmonic trap the relative strength $\epsilon_{{\rm dd}}=a_{{\rm dd}}/a_{s}$ between the DDI and the short-range interactions determines the many-body phase of the system Ferrier-Barbut et al. (2016). In particular, the dipolar length of 164Dy atoms is $a_{{\rm dd}}=\mu_{0}\mu^{2}_{m}m/12\pi\hbar^{2}=131a_{B}$, with $a_{B}$ being the Bohr radius. For sufficiently small values of $\epsilon_{{\rm dd}}$, the system exhibits a SF phase. The equilibrium solution is determined by the delicate balance between the attractive or repulsive long-range dipolar interaction and the repulsive contact interaction. As the short-range scattering length decreases, the relative strength of the long-range dipolar interaction becomes dominant. Within a small range of $\epsilon_{\rm dd}$, the system favors the SS phase, where a periodic structure of localized densities coherently connected by a dilute superfluid background emerges. However, a further increase in $\epsilon_{{\rm dd}}$ leads to a DL phase.
The initial state is prepared in an external 3D tilted double-well potential,
@@
where $V_{D}$ is the barrier height and $w$ is the barrier width. For our purposes, we use $V_{D}=10\hbar\omega_{x}$ and $w=0.5l_{{\rm osc}}$, with $l_{{\rm osc}}=\sqrt{\hbar/(m\omega_{x})}$ being the harmonic oscillator length and $\omega_{x}$ denoting the frequency of the external confinement. The last term is the linear external field gradient of strength $D$ that energetically favors the population on the left well. The tilt strength ranges from $D=1\hbar\omega_{x}/l_{{\rm osc}}$ to $D=10\hbar\omega_{x}/l_{{\rm osc}}$.
@@
Our analysis extends from (i) an elongated double-well characterized by $(\omega_{x},\omega_{y},\omega_{z})=2\pi\times(19,53,81)\rm Hz$ (with the single wells being similar to those in Tanzi et al. (2019b); Böttcher et al. (2019)) to (ii) a pancake double-well with $(\omega_{x},\omega_{y},\omega_{z})=2\pi\times(43,43,133)\rm Hz$ (as in Klaus et al. (2022)). The lowest panels of Fig. 1 show the contours of the potentials cut through $z=0$ for $D=10\hbar\omega_{x}/l_{\rm osc}$. Also, the upper panels (a)-(c) and (d)-(f) represent the dBEC configurations in the left well before the quench for the elongated and planar single wells respectively. The characteristic timescales determined by the trap frequency accordingly correspond to $\omega_{x}^{-1}=8.4$ ms and $\omega_{x}^{-1}=3.7$ ms in the elongated and planar cases, with harmonic oscillator lengths $l_{{\rm osc}}=\sqrt{\hbar/(m\omega_{x})}=1.8{\rm\mu m}$ and $1.2{\rm\mu m}$, respectively. Another important timescale is set by the energy difference among the pre- and post-quench states, $\tau_{D_{{\rm in}}}=1/\Delta\mathcal{E}_{D_{\rm in}}$, where $\Delta\mathcal{E}_{D_{\rm in}}\equiv\mathcal{E}_{D_{{\rm in}}}-\mathcal{E}_{D_{\rm fi}}$. Here, $\mathcal{E}_{D_{{\rm in}}}$ and $\mathcal{E}_{D_{{\rm fi}}}$ refer to the energy of the dBEC at the initial ($D_{{\rm in}}=D$) and final ($D_{{\mathrm{f}i}}=0$) tilt strength. The interaction dependence of $\Delta\mathcal{E}_{D_{\rm in}}$ is weak for the considered values of $\epsilon_{{\rm dd}}$. We thus provide an average value for each $D_{{\rm in}}$. In this sense, relevant timescales obtained from this energy difference correspond to $\tau_{D_{{\rm in}}=10}\approx 40$ms and $\tau_{D_{{\rm in}}=7}\approx 20$ms in both settings.
@@
In what follows, we explore a 164Dy dBEC under the influence of the energy offset for specific values of $\epsilon_{{\rm dd}}$ corresponding to a SF, a SS, and a DL in the initial state. After preparing the initial configuration, we perform a quench of the tilt strength from a finite value to zero, and monitor the system dynamics as the energy offset vanishes for up to one second.
@@
Figure 1: Ground state density iso-surfaces of a dBEC initially localized on the left side of (a)-(c) an elongated and (d)-(f) pancake-like tilted double-well trap, as depicted by the contour lines in the lowest panels. The density isosurfaces represent 20% and 2.5% of maximum density. All atoms reside in the left part of the double-well, and the entire density profile is located in the minimum of the left well. Depending on the relative interaction strength $\epsilon_{{\rm dd}}$ the dBEC in the left well features (a), (d) a regular SF for $\epsilon_{dd}=1.31$, (b) an elongated ($\epsilon_{{\rm dd}}=1.4$) or (e) planar SS ($\epsilon_{{\rm dd}}=1.49$), (c) an elongated DL pattern for $\epsilon_{dd}=1.49$ as well as (f) a single droplet surrounded by a homogeneous density distribution for $\epsilon_{{\rm dd}}=1.559$. Notice that the individual transitions are shifted to larger $\epsilon_{{\rm dd}}$ from the quasi-1D to the quasi-2D setting. The tilt strength is $D=7\hbar\omega_{x}/l_{{\rm osc}}$ enforcing all $N=40000$ magnetic atoms to reside on the left part of the double-well. The latter is characterized by frequencies (a)-(c) $(\omega_{x},\omega_{y},\omega_{z})=2\pi\times(19,53,81)\rm Hz$ and (d)-(f) $(\omega_{x},\omega_{y},\omega_{z})=2\pi\times(43,43,131)\rm Hz$, while having a width $w=0.5l_{{\rm osc}}$ and height $V_{D}=10\hbar\omega_{x}$. The potential contours at $z=0$ plane for $D=7\hbar\omega_{x}/l_{{\rm osc}}$, along with the variation of the potentials $V(x,0,0)$ across $x$ for $D=7\hbar\omega_{x}/l_{{\rm osc}}$ (solid line) and $D=0$ (dashed line), are shown in the left and right panels for elongated and pancake-like single-wells, respectively.